A tent of a given volume has a square base of side 2a, has its four sides vertical of height b and is surmounted by a regular pyramid of height 1. find the values of a and b in terms of height I such that the canvas required for its construction is minimum.

1
Expert's answer
2021-12-14T18:33:26-0500

area of canvas:

"S=4a\sqrt{h^2+a^2}-2ah"


"S'_h=\frac{4ha^2}{\sqrt{a^2h^2+a^4}}-2a=0"


"4h^2=h^2+a^2"

"a=h\sqrt 3"


"b^2=a^2+h^2=4h^2"

"b=2h"